### v1

# Scoring, Reasoning, and Selecting the Best! Ensembling Large Language Models via a Peer-Review Process

Abstract
We propose LLM-PeerReview, an unsupervised LLM Ensemble method that selects the most ideal response from multiple LLM-generated candidates for each query, harnessing the collective wisdom of multiple models with diverse strengths. LLM-PeerReview is built on a novel, peer-review-inspired framework that offers a clear and interpretable mechanism, while remaining fully unsupervised for flexible adaptability and generalization. Specifically, it operates in three stages: For scoring, we use the emerging LLM-as-a-Judge technique to evaluate each response by reusing multiple LLMs at hand; For reasoning, we can apply a principled graphical model-based truth inference algorithm or a straightforward averaging strategy to aggregate multiple scores to produce a final score for each response; Finally, the highest-scoring response is selected as the best ensemble output. LLM-PeerReview is conceptually simple and empirically powerful. The two variants of the proposed approach obtain strong results across four datasets, including outperforming the recent advanced model Smoothie-Global by 6.9% and 7.3% points, respectively.
Machine Learning, ICML
1 Introduction
The artificial intelligence domain has undergone a massive transformation recently, driven by the emergence of Large Language Models (LLMs) such as Gemini (Team et al., 2023), GPT-4 (Achiam et al., 2023), Llama (Touvron et al., 2023), and DeepSeek (Liu et al., 2024a). The success of these models has triggered a surge in research activity, with over 182,000 models now available on Hugging Face.
Behind this research enthusiasm, we can observe two main points (Chen et al., 2025; Jiang et al., 2023): 1) Persistent performance concerns: Although large language models can be easily deployed for zero-shot or in-context few-shot inference, they still face common performance issues, such as limited accuracy, hallucinations, and misalignment with human goals; 2) The varying strengths and weaknesses of LLMs: These models display significant behavioral differences, primarily driven by variations in their architecture, scale, training data, dictionary, tokenization and methodology. Consequently, their responses to the same prompt often diverge. With the above two points in mind and inspired by the spirit of Ensemble Learning (Dong et al., 2020), it is reasonable to suggest that relying on a single LLM—even one with a high public ranking or other criteria—may not be the optimal strategy for every user query. Instead, it might be more advantageous to simultaneously consider multiple LLM candidates (which are usable out-of-the-box) and leverage their distinct strengths. This concept is the core focus of the burgeoning field of LLM Ensemble (Chen et al., 2025).
As LLM Ensemble gains increasing attention, one well-established class of solutions—ensemble-after-inference (also known as post-hoc ensemble) methods (Jiang et al., 2023; Lv et al., 2024; Tekin et al., 2024; Li et al., 2024c; Guha et al., 2024; Si et al., 2023)—has emerged. These methods include the following two representative approaches (Chen et al., 2025):
  • Selection-then-regeneration approach (Jiang et al., 2023; Lv et al., 2024; Tekin et al., 2024), during inference on a downstream task, first employs a pre-trained “PairRanker” module to select the top-K candidate responses—those deemed most likely to be of high quality—from a pool of LLM-generated responses. This selected subset is then fed into another fine-tuned LLM (e.g., Flan-T5-XL (Chung et al., 2024)) to synthesize a final response. While this line of work has attracted significant attention (Jiang et al., 2023), they rely heavily on carefully curated task-specific training data and the need to fine-tune an additional LLM, limiting their generalization and adaptability.
  • Similarity-based selection approach (Li et al., 2024c; Guha et al., 2024; Si et al., 2023), instead, are mostly fully unsupervised (Guha et al., 2024; Si et al., 2023). These methods follow a simple and intuitive principle: for a given query, select the response with the highest total similarity to all other responses. While such methods pioneered unsupervised post-hoc LLM ensemble, their design remains coarse-grained—they rely on the naive similarity-based selection strategy (Li et al., 2024c; Guha et al., 2024), along with shallow similarity measure of BLEU (Li et al., 2024c) or limited informational utilization (Guha et al., 2024). Thus, the true potential of selection-based post-hoc ensemble remains largely untapped.
When we revisit this research problem, we ask the most fundamental question: In the real world, how would humans select the most ideal text from a set of candidate texts? Perhaps the most immediate and relatable real-world example is: the academic peer-review process.
Motivated by this, we propose a new, fully unsupervised LLM Ensemble method called LLM-PeerReview. Specifically, LLM-PeerReview is structured sequentially around three components: 1) Scoring (analogous to paper reviewing): Given multiple candidate responses to the same query, we adopt the LLM-as-a-Judge approach—leveraging available LLMs as evaluators that assess each response and assign a score (e.g., 5.0 indicating Strong Accept, 4.0 indicating Weak Accept, etc.). To improve scoring accuracy and reduce bias, we further introduce a technique termed “flipped-triple scoring trick” into the Scoring process; 2) Reasoning (analogous to final score estimation made by the senior reviewer): Besides being able to perform direct averaging calculations, we can invoke graphical model-based truth inference techniques from crowdsourcing and weak supervision literature, to perform refined, reliability-aware weighted score aggregation, deriving a final score for each response; 3) Selection (analogous to final decision made by the senior reviewer): This step is analogous to how a senior reviewer or area chair selects the most suitable paper from a small set of submissions. For each query, once final scores have been inferred for all responses, we select the highest-scoring response as the ensemble result.
LLM-PeerReview is built upon the unsupervised, selection-based paradigm, and introduces a novel peer-review-inspired framework for LLM Ensemble, offering a clear and interpretable mechanism. Within the methodology, LLM-PeerReview leverages the emerging LLM-as-a-Judge technique (Li et al., 2024a, b) to evaluate each candidate response, thereby effectively reusing the collective intelligence of multiple existing LLMs at hand. Moreover, the use of a graphical-model-based truth inference algorithm allows us to benefit from the principled graphical model for refined and reliability-aware aggregation of multiple scoring signals. Empirically, extensive experiments conducted with 7B-scale LLMs show that the proposed LLM-PeerReview outperforms recent advanced similarity-based methods (Li et al., 2024c; Guha et al., 2024).
1.1 Related Work
LLM Ensemble, as outlined in Section 1, can be broadly categorized into three approaches (Chen et al., 2025). The first category, ensemble-before-inference approach (Shnitzer et al., 2023; Srivatsa et al., 2024; Ong et al., 2024), typically necessitates custom-labeled data to pretrain a classifier that routes each query. The mandatory pretraining phase and the dependency on labeled data are the primary inconveniences of these methods. The second category, ensemble-during-inference approach, can be subdivided into token-level (Yu et al., 2024; Huang et al., 2024; Xu et al., 2024), span-level (Liu et al., 2024b; Xu et al., 2025), and process-level (Park et al., 2024) methods, depending on the level of granularity of the information considered in the ensemble. These methods, however, entail substantial computational costs and require the local deployment of every LLM in the ensemble. Lastly, as mentioned before, ensemble-after-inference approach mainly includes selection-then-regeneration-based (Lu et al., 2024) and aggregation-based methods (Li et al., 2024c; Guha et al., 2024).
Figure 1: The proposed LLM-PeerReview contains three steps: (1) Scoring: For a given query, after each LLM independently generates a response (analogous to a submitted academic paper), LLM-PeerReview applies the LLM-as-a-Judge technique (and the proposed flipped-triple scoring trick), treating each model as a reviewer to assign scores to all candidate responses; (2) Reasoning: LLM-PeerReview then uses a truth inference algorithm—analogous to a senior reviewer—to estimate a final score for each response. (Notably, for the variant LLM-PeerReview-Weighted, the inference algorithm is performed using score information across all queries, allowing the model to learn each LLM’s scoring behavior using global information from the dataset, thereby enabling fine-grained, reliability-aware score aggregation); (3) Selecting the best: Finally, for each query, LLM-PeerReview selects the response with the highest final score as the ensemble output—analogous to how a senior reviewer chooses the best paper from a specific submission pool.
Weak Supervision (Zhang et al., 2021; Chen et al., 2023b), also commonly referred to as Learning from Crowds (Chen et al., 2021b, 2022), is a research problem that closely resembles post-inference ensemble methods. The key distinction is that Weak Supervision methods either focus on learning classifiers directly from imperfectly labeled data (Zhang et al., 2021) or on aggregating weak label information (Zheng et al., 2017; Chen et al., 2022), whereas ensemble-after-inference LLM Ensemble methods are exclusively concerned with aggregation and do not involve the learning of classifier models. Additionally, a significant difference is that most weakly supervised learning methods are primarily focused on classification scenarios with closed answer sets, rather than text generation tasks that involve open-ended output spaces.
LLM-as-a-Judge approaches has received a lot of attention recently. As task complexity increases and model outputs become more diverse, traditional evaluation methods—such as matching-based or embedding-based metrics—often fail to capture subtle attributes and provide reliable results (Gu et al., 2024). The recent emergence of large language models has led to the development of the “LLM-as-a-judge” paradigm, where LLMs assess the quality of model outputs. These methods can be broadly classified into three categories: Single-LLM-based, Multi-LLM-based, and Human-AI collaboration-based evaluation approaches. Single-LLM-based methods primarily focus on prompt design (Fu et al., 2023; Kotonya et al., 2023), fine-tuning (Chen et al., 2023a; Wu et al., 2024), or post-processing (Daynauth & Mars, 2024). Notably, Multi-LLM-based methods, which include collaborative (Zhang et al., 2023), competitive (Owens et al., 2024), and aggregation-based (Verga et al., 2024; Chen et al., 2024; Chu et al., 2024) strategies, are closely related to our approach (especially for aggregation methods).
2 LLM-PeerReview
This section presents our proposed method, LLM-PeerReview, with an overview shown in Figure 1. We begin by formalizing the research problem, followed by a detailed introduciton of the three components of our method in Sections 2.12.2, and 2.3.
Problem Formulation: (Unsupervised) LLM Ensemble.
Without access to any reference responses (i.e., ideal/truth responses), we are given a set of queries $\{\mathbf{x}^{(i)}\}_{i=1}^{I}$ for a generative task. We have access to $J$ large language models $\{\mathcal{M}_{j}\}_{j=1}^{J}$, where each model $\mathcal{M}_{j}$ generates a response $\mathbf{r}^{(i,j)}=\mathcal{M}_{j}(\mathbf{x}^{(i)})$—which is often not ideal—for a given query $\mathbf{x}^{(i)}$. Thus, for each query, we have a set of zero-shot inference responses $\mathbf{R}^{(i)}=[\mathbf{r}^{(i,1)},\ldots,\mathbf{r}^{(i,J)}]$ from heterogeneous LLMs $\{\mathcal{M}_{j}\}_{j=1}^{J}$, while the underlying reference response $\mathbf{y}^{(i)}$ is unobserved to us. Our goal is to ensemble the LLM responses to produce a single, high-quality final response for each query $\mathbf{x}^{(i)}$, using the available data $\mathcal{D}=\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\}_{i=1}^{I}$.
2.1 Scoring
Naive point-wise scoring.
As shown in Figure 1 and within our proposed LLM-PeerReview, the scoring phase occurs after the LLMs have first generated responses to the input queries. Using the LLM-as-a-Judge technique (Gu et al., 2024), each LLM judge can assign a point-wise score to each response, representing its overall quality. For example, the score can range from [1, 2, 3, 4, 5], representing the levels of [“Very Poor”, “Poor”, “Acceptable”, “Good”, “Excellent”]. 1111) It is worth noting that each scoring prompt contains both the corresponding query and the response to be evaluated, rather than presenting the response alone; 2) The scoring prompts that we designed and utilized in the experiments are provided in the Appendix.
Flipped-triple scoring trick.
The above naive point-wise scoring technique can provide scores; however, we propose a technique called flipped triple scoring, which we recommend when applying our approach. Specifically: 1) For multiple responses from different models to the same query $\mathbf{x}^{(i)}$, we first shuffle them; 2) Then, for each LLM judge $\mathcal{M}_{j^{\prime}}$, we score the response triplet $\mathbf{r}^{(i,j-1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j+1)}]$ sequentially (with $J$ times), and for each iteration, we also score the flipped version (i.e., $\mathbf{r}^{(i,j+1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j-1)}]$). As a result, each response receives six scores from the same LLM judge. We can simply average these scores to obtain a final score, which serves as the score $y^{(i,j;j^{\prime})}$ for response $\mathbf{r}^{(i,j)}$ by LLM judge $\mathcal{M}_{j^{\prime}}$. In short, this technique mitigates two common scoring biases in LLM-as-a-Judge (Wang et al., 2023; Zheng et al., 2023; Gu et al., 2024). First, in point-wise scoring, models tend to show a consistent bias toward certain scores (e.g., consistently assigning a score of “1”), as they evaluate a single response without the reference effect of multiple responses. Also, when multiple responses (such as two or three) are presented for evaluation at one time, models frequently exhibit a position bias, tending to favor responses that appear either at the beginning or at the end.
2.2 Reasoning: a Truth Inference Process
First variant: LLM-PeerReview-Average.
After the scoring phase, as shown in Figure 1, each response $\mathbf{r}^{(i,j)}$ corresponding to a query $\mathbf{x}^{(i)}$ receives multiple scores $\{y^{(i,j;j^{\prime})}\}_{j^{\prime}=1}^{J}$ provided by multiple LLM judges. Then, how can we aggregate these scores meaningfully to compute a final, reliable score $\hat{t}^{(i,j)}$ for each response $\mathbf{r}^{(i,j)}$? This problem is analogous to designing an algorithm that simulates a senior reviewer who consolidates evaluations from multiple reviewers with different scoring preferences and evaluation capabilities. A straightforward and intuitive approach is “averaging” (Zheng et al., 2017; Zhang et al., 2021)—simply taking the mean of all the scores for a given response. As a simple variant of our approach, we refer to it as LLM-PeerReview-Average.
Second variant: LLM-PeerReview-Weighted.
We observe that the above averaging strategy assumes all models to be equally reliable, ignoring the inherent differences in evaluation quality across models. Considering this point, we further propose a weighted variant, referred to as LLM-PeerReview-Weighted. Here, we invoke the well-established Dawid-Skene (DS) model (Dawid & Skene, 1979), a canonical truth-inference graphical model widely used in weak supervision learning and crowdsourcing (Zheng et al., 2017; Zhang et al., 2021), and adapt it to our context. In the following, we introduce the construction of the graphical model in Section 2.2.1, and present the optimization objective and optimization (to obtain the final score value for each response) in Section 2.3.
2.2.1 Graphical Model
Figure 2: Probabilistic graphical representation.
Overall, to infer the underlying “truth” score (unobserved) behind the multiple weak/non-ideal score annotations (observed) for each response, we construct a latent variable graphical model (Bishop & Nasrabadi, 2006; Everett, 2013; Chen et al., 2023b) that includes a latent variable representing the truth score. As depicted the graphical representation in Figure 2, we next introduce the probabilistic generative process we construct from truth scores to weak scores labeled by LLM judges.
First, for each response $\mathbf{r}^{(i,j)}$, we assume that its true score $t^{(i,j)}$ is drawn from a categorical distribution:
$t^{(i,j)}\sim\operatorname{Cat}(t^{(i,j)};\bm{\alpha})\mathrm{,}$ (1)
where the distribution is parametrized by $\bm{\alpha}$. Next, similar to the concept of confusion matrix commonly used in machine learning (Bishop & Nasrabadi, 2006; Goodfellow, 2016), we introduce an annotator-specific transition matrix $\bm{\Pi}^{(j^{\prime})}$ to model the probability that an LLM confuses one score category for another, capturing its scoring tendencies and potential biases:
$p(y^{(i,j;j^{\prime})}=n|t^{(i,j)}=m;\bm{\Pi}^{(j^{\prime})})=\pi_{mn}^{(j^{\prime})}\mathrm{,}$ (2)
where $m,n\in\{1,\ldots,K\}$ and $K$ denotes the number of categories (i.e., the number of score levels).
2.2.2 Objective and Optimization
Objective.
Based on the model construction above, the optimization objective is to maximize the log conditional likelihood of the observed scoring labels $\mathbf{Y}=\{y^{(i,j;j^{\prime})}\mid 1\leq i\leq I,\ 1\leq j\leq J,\ 1\leq j^{\prime}\leq J\}$ contributed by $J$ LLM judges, i.e., $\log p(\mathbf{Y};\Theta)$ w.r.t. the parameters $\Theta=\{\bm{\alpha},\bm{\Pi}^{(1)},\ldots,\bm{\Pi}^{(J)}\}$.
Optimization.
In brief, as with most latent variable models (Bishop & Nasrabadi, 2006; Dawid & Skene, 1979), we apply the Expectation-Maximization (EM) algorithm (Dempter, 1977) to solve the optimization problem. [^2] We provide the detailed derivations in Appendix. First, the log-likelihood can be written as:
$\log p(\mathbf{Y};\Theta)=\sum_{i=1}^{I}\sum_{j=1}^{J}\log p(\mathbf{y}^{(i,j)};\Theta)$ (3)
where $\mathbf{y}^{(i,j)}=\{y^{(i,j;j^{\prime})}\mid 1\leq j^{\prime}\leq J\}$ denotes the set of scores assigned to response $\mathbf{r}^{(i,j)}$ by the $J$ models. Then, we use Jensen’s inequality (Bishop & Nasrabadi, 2006) to derive the Evidence Lower Bound (ELBO):
$\log p(\mathbf{Y};\Theta)\geq\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{t^{(i,j)}}q(t^{(i,j)})\log\frac{p(\mathbf{y}^{(i,j)},t^{(i,j)};\Theta)}{q(t^{(i,j)})},$ (4)
where $q(t^{(i,j)})$ is a discrete distribution over the variable $t^{(i,j)}$. In the following, we then proceed to apply the general EM recipe to perform iterative calculations (concerning E-step and M-step) to solve the optimization problem $\Theta:=\underset{\Theta}{\operatorname{argmax}}\log p(\mathbf{Y};\Theta)$.
E-step (inference).
The posterior $q(t^{(i,j)})$ is obtained by using of Bayes’s theorem given the parameters $\Theta=\{\bm{\alpha},\bm{\Pi}^{(1)},\ldots,\bm{\Pi}^{(J)}\}$ learned on the last M-step:
$q(t^{(i,j)}=k)=p(t^{(i,j)}=k\mid\mathbf{y}^{(i,j)};\Theta)$ (5)
$\propto p(t^{(i,j)}=k;\bm{\alpha})\cdot\prod_{j^{\prime}=1}^{J}p(y^{(i,j;j^{\prime})}\mid t^{(i,j)}=k;\bm{\Pi}^{(j^{\prime})}).$
Given that we are likely to obtain decimal values rather than integers on $y^{(i,j;j^{\prime})}$ after the scoring phase, we make the adaptation:
$q(t^{(i,j)}=k)\propto$ $p(t^{(i,j)}=k;\bm{\alpha})\cdot\prod_{j^{\prime}=1}^{J}[\phi_{l}\cdot p(y^{(i,j;j^{\prime})}_{l}\mid t^{(i,j)}=k;$ (6)
$\bm{\Pi}^{(j^{\prime})})$ $+\phi_{u}\cdot p(y^{(i,j;j^{\prime})}_{u}\mid t^{(i,j)}=k;\bm{\Pi}^{(j^{\prime})})],$
[…]
M-step (learning).
Furthermore, by maximizing optimization objective in Equation 4 and using the standard Lagrange multiplier method (Bishop & Nasrabadi, 2006), we can obtain the closed-form solution for $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ in Equation 7 shown below; and by equating the gradient of Equation 4 to zero, we can obtain the closed-form solution for $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ in Equation 8 shown below.
$\alpha_{k}=\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=k)}{I\cdot J},$ (7)
$\pi_{mn}^{(j^{\prime})}=\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=m)\cdot\Psi(y^{(i,j;j^{\prime})},n)}{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=m)}\mathrm{,}$ (8)
where $\Psi(y^{(i,j;j^{\prime})})=[\phi_{l}\cdot\mathbb{I}(y^{(i,j;j^{\prime})}_{l}=n)+\phi_{u}\cdot\mathbb{I}(y^{(i,j;j^{\prime})}_{u}=n]$, and $\mathbb{I}(\cdot)$ is an indicator function that takes the value $1$ when the internal declaration is true, and 0 otherwise.
Obtain the final score value for each response.
After the EM-based optimization, we obtain the posterior probabilities $q(t^{(i,j)})$ over score categories for each response. Then, we compute the final score for each response through the following simple summation:
$S(t^{(i,j)})=\sum_{k=1}^{K}q(t^{(i,j)}=k)\cdot s_{k},$ (9)
where $s_{k}$ denotes the score value corresponding to the $k$-th scoring category. [^3] For example, suppose that for a given response, we have $q(t^{(i,j)}=4)=0.5$, $q(t^{(i,j)}=5)=0.5$, along with the score values $s_{4}=4.0$, $s_{5}=5.0$. Then, the final aggregated score is $S(t^{(i,j)})=0.5\cdot 4.0+0.5\cdot 5.0=4.5$.
2.3 Selecting the Best
Finally, for each query $\mathbf{x}^{(i,j)}$, we can easily determine its optimal response:
$\mathbf{r}_{\text{ensemble}}^{(i)}=\underset{\mathbf{r}^{(i,j)}}{\operatorname{argmax}}\{S(t^{(i,j)})\mid 1\leq j\leq J\},$ (10)
which is selected as the final result after the ensemble.
The overall procedure for our proposed LLM-PeerReview is summarized in Algorithm 1.
Algorithm 1 LLM-PeerReview-Weighted
Input: Data $\mathcal{D}=\left\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\right\}_{i=1}^{I}$, where for each query $\mathbf{x}^{(i)}$, we have responses $\mathbf{R}^{(i)}=\{\mathbf{r}^{(i,j)}\mid 1\leq j\leq J\}$ from heterogeneous LLMs $\left\{\mathcal{M}_{j}\right\}_{j=1}^{J}$
Output: results $\{\mathbf{r}_{\text{ensemble}}^{(i)}\}_{i=1}^{I}$
1: # 1) Scoring:
2: Each LLM $\mathcal{M}_{j^{\prime}}$ acts as a judge and assigns a score $y^{(i,j;j^{\prime})}$ to each response $\mathbf{r}^{(i,j)}$
3:# 2) Reasoning:
4:# 2.1) Reasoning the posterior probabilities $q(t^{(i,j)})$ over score categories for each response:
5: Initialize posterior $\{q(t^{(i,j)})\mid 1\leq i\leq I,1\leq j\leq J\}$ by averaging scores on $\textbf{r}^{(i,j)}$
6:while not converge do
7:  Update $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ by using Eq. 7
8:  Update $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ by using Eq. 8
9:  Update the posterior $\{q(t^{(i,j)})\mid 1\leq i\leq I,1\leq j\leq J\}$ by using Eq. 10
10:end while
11:# 2.2) Obtain the final score value for each response:
12: Obtain the final score value $S(t^{(i,j)})$ for each individual response $\mathbf{r}^{(i,j)}$ by using Eq. 9
13:# 3) Selecting the best:
14: Obtain the final results $\{\mathbf{r}_{\text{ensemble}}^{(i)}\}_{i=1}^{I}$ by using Eq. 10
[…]
Single LLM Llama-3.1-8B-Instruct 75.3 79.3 52.3 7.3 53.5
Mistral-7B-Instruct 72.7 64.3 26.5 10.4 43.5
Qwen2-7B-Instruct 63.0 88.5 59.8 15.2 56.6
Qwen2.5-7B-Instruct 62.5 91.5 69.3 27.6 62.7
Theoretical average 68.4 80.9 51.9 15.1 54.1
LLM Ensemble Random 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global 63.0 91.5 59.8 27.6 60.5
Smoothie-Local 73.6 85.5 61.8 18.3 59.8
Agent-Forest 70.5 86.8 61.0 22.1 60.1
GaC 71.5 91.8 54.0 23.6 60.2
LLM-PeerReview-Average 76.9 $\pm$ 0.1 92.7 $\pm$ 0.3 69.5 $\pm$ 0.2 30.4 $\pm$ 0.1 67.4
LLM-PeerReview-Weighted 77.0 $\pm$ 0.1 93.0 $\pm$ 0.2 71.0 $\pm$ 0.2 30.2 $\pm$ 0.1 67.8
Our variants Llama-3-8B-Selection 76.5 $\pm$ 0.2 90.8 $\pm$ 0.6 68.8 $\pm$ 0.5 29.6 $\pm$ 0.3 66.4
Mistral-7B-Selection 75.6 $\pm$ 0.3 90.8 $\pm$ 0.1 66.4 $\pm$ 0.3 25.9 $\pm$ 0.4 64.7
Qwen2-7B-Selection 74.2 $\pm$ 0.2 88.8 $\pm$ 0.6 61.7 $\pm$ 0.7 23.7 $\pm$ 0.3 62.1
Qwen2.5-7B-Selection 75.5 $\pm$ 0.2 92.1 $\pm$ 0.4 66.2 $\pm$ 0.6 28.1 $\pm$ 0.1 65.5
Table 1: Main results (%).
Figure 3: LLM performances (bottom: AlpacaEval).
3 Experiments
3.1 Setup
We provide more details in Appendix A, including experimental setups, additional results, and the prompts used.
Datasets and evaluation.
We evaluate four widely-used datasets  [^4] We employ the same versions as those used in previous studies (Guha et al., 2024; Hu et al., 2024)., grouped into three categories: (1) Factual Recall: TriviaQA (Joshi et al., 2017; Guha et al., 2024) evaluates the accuracy of model responses to factual questions across various domains, including history, science, and geography. (2) Arithmetic Reasoning: GSM8k (Chen et al., 2021a; Hu et al., 2024) and MATH (Hendrycks et al., 2021; Hu et al., 2024) assess basic arithmetic and more advanced mathematical reasoning, respectively, with accuracy as the evaluation metric, focusing on correct numerical answers. (3) Instruction Following: AlpacaEval (Dubois et al., 2023; Hu et al., 2024) tests models’ ability to follow various instructions. We use GPT-4o-mini to evaluate the accuracy of model responses, assessing whether the model’s response exceeds the reference answer in the dataset.
Seed LLMs for ensemble.
Considering 7B-scale models are widely used by researchers and generally regarded as having acceptable judging capabilities (Wang et al., 2025; Kim et al., 2024), we use these well-established 7B models for ensemble: Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. Furthermore, we employ the instruction-tuned versions of these models rather than their base counterparts. This choice allows us to leverage their instruction-following capabilities, which are essential for the LLM-as-a-judge tasks.
Baselines.
We compare the proposed LLM-PeerReview with the two categories of baselines. (1) Single LLMs: The four 7B-scale models, Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. (2) LLM Ensemble baselines: (i) Random (Lu et al., 2024; Guha et al., 2024) is a random-selection baseline that simply returns the response from a randomly chosen LLM in the ensemble. As one of the simplest ensemble strategies for large language models, this method has previously been applied to dialogue tasks (Lu et al., 2024); (ii) Smoothie-Global (Guha et al., 2024), Smoothie-Local (Guha et al., 2024), and Agent-Forest (Li et al., 2024c) are recently proposed, strong similarity-based ensemble methods, as introduced in detail in Section 1; (iii) GaC (Yu et al., 2024) is a representative token-level ensemble-during-inference approach. It constructs a unified vocabulary that merges the individual dictionaries of multiple LLMs. During inference, token sampling is performed by observing the output distributions from these models across the unified vocabulary.
Configurations.
(1) For each individual large language model, we follow the setup of Smoothie (Guha et al., 2024), where the model responds once to each query. The responses from all models are stored for integration by the LLM Ensemble methods. (2) For the two variants of the baseline Smoothie (Guha et al., 2024), we set the number of neighbors as specified in the original paper. Agent-Forest (Li et al., 2024c) does not require any hyperparameter configuration. For our method, we set the model temperature to 0 during the scoring process to eliminate suboptimal results caused by randomness. Additionally, the scoring prompts used across the four datasets are provided in the Appendix. (3) All experiments were performed using 6 or 4 parallel Nvidia V100 32GB GPUs. All experiments with stochastic outputs were conducted three times.
(a) Left: The transition matrix of each LLM estimated by LLM-PeerReview-Weighted. Right: Correlation between matrix diagonal information of each LLM and its performance as a single judge (corresponding to “our variants” in Table 1). For the first three datasets with ground-truth answers, diagonal information is represented by extreme values $(\pi_{11}^{(j^{\prime})}+\pi_{KK}^{(j^{\prime})})$; for the instruction-following dataset AlpacaEval, the sum of all diagonal values is used.
Method TriviaQA GSM8k MATH AlpacaEval Average
Variant Performance ($\uparrow$):
   Random 65.7 $\pm$ 1.3 81.3 $\pm$ 0.6 51.2 $\pm$ 1.8 14.7 $\pm$ 0.5 53.2
   Single 69.2 $\pm$ 0.6 85.5 $\pm$ 2.1 60.3 $\pm$ 1.6 23.8 $\pm$ 1.0 59.7
   Double 73.3 $\pm$ 0.5 90.0 $\pm$ 0.7 71.3 $\pm$ 0.2 29.2 $\pm$ 0.2 66.0
   Flipped-triple 74.5 $\pm$ 0.0 90.8 $\pm$ 0.2 71.5 $\pm$ 0.4 30.5 $\pm$ 0.0 66.8
   Quadruple-half 74.7 $\pm$ 0.2 91.5 $\pm$ 0.4 73.3 $\pm$ 0.2 29.2 $\pm$ 0.2 67.2
Computation Efficiency (seconds/scoring the 4 responses for each sample; $\downarrow$):
   Single ($\mathcal{O}(J)$) 7.89 10.2 10.6 16.9 11.4
   Double ($\mathcal{O}(J^{2})$) 37.1 49.4 51.6 77.4 53.9
   Flipped-triple ($\mathcal{O}(J)$) 29.7 43.4 47.1 74.3 48.6
   Quadruple-half ($\mathcal{O}(J!)$) 51.3 83.8 90.0 137.65 90.7
Table 2: Top: Performance of the base variant LLM-PeerReview-Average with different scoring strategies. Note that, for computation efficiency, we use 200 samples from each dataset (all shuffled as necessary). (i) Random refers to the same baseline in Table 1; (ii) Single refers to scoring a single response at a time; (iii) Double refers to scoring all response pairs within the response set at a time; (iv) Quadruple-half refers to scoring all possible response quadruple sequences in the response set; given the high computational cost, we used a relaxed version and only calculated half of the cases. Bottom: Computation efficiency.
Figure 11: Performance of various variants across different scoring levels.
3.2 Main Results
The ensemble of the proposed LLM-PeerReview is effective. The main results are shown in Table 1. First, by examining the results in the “Single LLM” and “LLM Ensemble” sections of Table 1, one key finding is that both of our variants consistently outperform any single LLM and all LLM Ensemble baselines across all datasets. In the last column, which presents the average performance, our two variant methods (with results of 67.4% and 67.8%) surpass the strongest single model, Qwen2.5, by 4.7% and 5.1%, respectively, and outperform the strongest ensemble method, Smoothie-Global, by 6.9% and 7.3%. These results directly demonstrate the effectiveness of our method, as it achieves superior performance by integrating the collective knowledge of multiple models across factual-recall QA tasks, mathematical reasoning tasks, and instruction-following tasks. Additionally, the ensemble task across these four datasets is challenging, as the performance of the four LLMs varies significantly for each dataset. In contrast, ensembling four LLMs with similar performance would make it easier to achieve superior results compared to any single LLM.
Each LLM has its strengths and weaknesses. In Figure 3, the upper subplot presents a radar chart of individual LLM performance, while the lower subplot displays the win-tie-loss chart for models on the challenging instruction-following dataset, AlpacaEval. This chart highlights that models with the best overall performance may underperform on specific tasks compared to those with weaker overall results. In summary, the results in Table 1 and Figure 3 demonstrate that a strong LLM does not excel across all datasets. Each model has its strengths and weaknesses, highlighting the substantial practical significance of LLM Ensemble.
3.3 Significance of Aggregating Multiple Judges
Simply averaging the scores from multiple judges is also quite effective. In the “Our variants” of Table 1, we present the performance of using a single LLM as a judge to select the optimal response. From the average performances in the last column of Table 1, we observe that these variants perform quite well (surpassing the overall best model, Qwen2.5, in 3/4 cases). However, when comparing the performance of these variants with that of our prototype LLM-PeerReview-Average, it becomes clear that aggregating and averaging the scores from multiple judges is highly beneficial, compared to relying solely on the score of a single large model.
The weighted truth inference has the potential for further performance improvement. By observing the average results in Table 1, we find that LLM-PeerReview-Weighted leads to further performance gains compared to simple averaging. In the left subplot of Figure 11(a), we observe subtle variations in the transition matrices learned for each model. On the other hand, the right subplot in Figure 11(a), displaying positive correlation coefficients, demonstrates that our method can effectively identifies stronger and weaker judges.
3.4 Alternative Designs
The flipped-triple scoring trick represents a performance-efficiency trade-off. As introduced in Section 2.1, it is intuitive that, in addition to the our recommended flipped-triple scoring method, several variant scoring methods could be employed (their definitions are provided in the caption of Table 2). Overall, the performance of these four variants follows the order: $\textit{quadruple-half}>\textit{flipped-triple}>\textit{double}>\textit{single}$. Variants quadruple-half, flipped-triple, and double all offer noticeable de-biasing performance advantages over the single-scoring strategy. On the other hand, in terms of theoretical computational complexity, the complexities for de-biased strategies double/flipped-triple/quadruple-half are $\mathcal{O}(J)$$\mathcal{O}(J^{2})$$\mathcal{O}(J)$$\mathcal{O}(J!)$, with strategy flipped-triple having the lowest computational complexity. Furthermore, in Table 2, we present the scoring efficiency of these four strategies. Compared to strategies double and quadruple-half, strategy flipped-triple is the most time-efficient.
Common scoring levels can generally be attempted. In Figure 11, we conduct a further analysis of how different scoring levels influence the performance of both our method and the four individual scoring models. [^5] As indicated by the prompts in the Appendix, for the main experiment in Table 1, the scoring levels applied across the four datasets were 5, 3, 3, and 10, respectively. For each scoring level, we have carefully crafted meaningful descriptions and corresponding prompts. We use the basic variant LLM-PeerReview-Avarage for this analysis. Under these conditions, our method exhibits slightly varying performance, showing no consistent tendencies across the levels of 3, 5, 7, and 10.
4 Conclusion
This paper presents LLM-PeerReview, an unsupervised peer-review-inspired method, to address the LLM ensemble problem. PeerReview benefits both from the analytical capabilities of the powerful large language models at hand when evaluating response quality, and from the straightforward averaging strategy or the principled, fine-grained score aggregation when inferring final scores. LLM-PeerReview—embedded with the well-established techniques of LLM-as-a-Judge—closely emulates the real-world human process of selecting the best text, offering a clear and interpretable mechanism. Our empirical evaluations on four datatests demonstrate that LLM-PeerReview significantly improves the recent advanced model Smoothie-Global and provides a new solution to LLM Ensemble.
Appendix A Appendix
In this appendix, we present the following sections to further support the content of the main paper.
  • Section A.1: The optimization process of the graphical model constructed in the main text;
  • Section A.2: More details on the experimental setup;
  • Section A.3: More experimental results;
  • Section A.4: The prompts used in the experiments.
A.1 Proof of the Optimization
The optimization objective is to maximize the log conditional likelihood of the observed scoring labels $\mathbf{Y}=\{y^{(i,j;j^{\prime})}\mid 1\leq i\leq I,\ 1\leq j\leq J,\ 1\leq j^{\prime}\leq J\}$ contributed by $J$ LLM judges, i.e., $\log p(\mathbf{Y};\Theta)$. In brief, as with most latent variable models (Bishop & Nasrabadi, 2006; Dawid & Skene, 1979), we apply the Expectation-Maximization (EM) algorithm (Dempter, 1977) to solve the optimization problem. First, the log-likelihood is:
$\log p(\mathbf{Y};\Theta)=$ $\log\prod_{i=1}^{I}\prod_{j=1}^{J}p(\mathbf{y}^{(i,j)};\Theta)$ (A.1)
$=$ $\sum_{i=1}^{I}\sum_{j=1}^{J}\log p(\mathbf{y}^{(i,j)};\Theta)$
$=$ $\sum_{i=1}^{I}\sum_{j=1}^{J}\log\sum_{t^{(i,j)}}p(\mathbf{y}^{(i,j)},t^{(i,j)};\Theta),$
$=$ $\sum_{i=1}^{I}\sum_{j=1}^{J}\log[q(t^{(i,j)})\sum_{t^{(i,j)}}\frac{p(\mathbf{y}^{(i,j)},t^{(i,j)};\Theta)}{q(t^{(i,j)})}],$
$\geq$ $\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{t^{(i,j)}}q(t^{(i,j)})\log\frac{p(\mathbf{y}^{(i,j)},t^{(i,j)};\Theta)}{q(t^{(i,j)})}$
$\geq$ $\sum_{i=1}^{I}\sum_{j=1}^{J}[\mathbb{E}_{q(t^{(i,j)})}\operatorname{log}p(t^{(i,j)};\bm{\alpha})+\mathbb{E}_{q(t^{(i,j)})}\log\prod_{j^{\prime}=1}^{J}p({y}^{(i,j;j^{\prime})}\mid t^{(i,j)};\bm{\Pi}^{(j^{\prime})})-\mathbb{E}_{q(t^{(i,j)})}\log q({t}^{(i,j)})],$
where $\mathbf{y}^{(i,j)}=\{y^{(i,j;j^{\prime})}\mid 1\leq j^{\prime}\leq J\}$ denotes the set of scores assigned to response $\mathbf{r}^{(i,j)}$ by the $J$ models. The derivation which obtains the Evidence Lower Bound (ELBO)
$\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{t^{(i,j)}}q(t^{(i,j)})\log\frac{p(\mathbf{y}^{(i,j)},t^{(i,j)};\Theta)}{q(t^{(i,j)})}$ (A.2)
used Jensen’s inequality (Bishop & Nasrabadi, 2006); and $q(t^{(i,j)})$ is a discrete distribution over the variable $t^{(i,j)}$. In the following, we then proceed to apply the general EM recipe to perform iterative calculations (concerning E-step and M-step) to solve the optimization problem: $\Theta:=\underset{\Theta}{\operatorname{argmax}}\log p(\mathbf{Y};\Theta)$.
$\textbf{E-step (inference):}\qquad q(t^{(i,j)}=k)$ $=p(t^{(i,j)}=k\mid\mathbf{y}^{(i,j)};\Theta)$ (A.3)
$\propto p(t^{(i,j)}=k;\bm{\alpha})\cdot\prod_{j^{\prime}=1}^{J}p(y^{(i,j;j^{\prime})}\mid t^{(i,j)}=k;\bm{\Pi}^{(j^{\prime})}).$
The posterior $q(t^{(i,j)})$ is obtained by using of Bayes’s theorem given the parameters $\Theta=\{\bm{\alpha},\bm{\Pi}^{(1)},\ldots,\bm{\Pi}^{(J)}\}$ learned on the last M-step. Given that we are likely to obtain decimal values rather than integers on $y^{(i,j;j^{\prime})}$ after the scoring phase, we make the adaptation:
$q(t^{(i,j)}=k)\propto p(t^{(i,j)}=k;\bm{\alpha})\cdot\prod_{j^{\prime}=1}^{J}[\phi_{l}\cdot p(y^{(i,j;j^{\prime})}_{l}\mid t^{(i,j)}=k;\bm{\Pi}^{(j^{\prime})})+\phi_{u}\cdot p(y^{(i,j;j^{\prime})}_{u}\mid t^{(i,j)}=k;\bm{\Pi}^{(j^{\prime})})],$ (A.4)
[…]
M- step (learning): (A.5)
$\Theta$= $\underset{\Theta}{\operatorname{argmax}}\sum_{i=1}^{I}\sum_{j=1}^{J}[\mathbb{E}_{q(t^{(i,j)})}\operatorname{log}p(t^{(i,j)};\bm{\alpha})+\mathbb{E}_{q(t^{(i,j)})}\log\prod_{j^{\prime}=1}^{J}p({y}^{(i,j;j^{\prime})}\mid t^{(i,j)};\bm{\Pi}^{(j^{\prime})})-\mathbb{E}_{q(t^{(i,j)})}\log q({t}^{(i,j)})]$
= $\underset{\Theta}{\operatorname{argmax}}\sum_{i=1}^{I}\sum_{j=1}^{J}[\mathbb{E}_{q(t^{(i,j)})}\operatorname{log}p(t^{(i,j)};\bm{\alpha})+\mathbb{E}_{q(t^{(i,j)})}\log\prod_{j^{\prime}=1}^{J}p({y}^{(i,j;j^{\prime})}\mid t^{(i,j)};\bm{\Pi}^{(j^{\prime})})].$
Furthermore, by maximizing optimization objective in Equation A.5 and using the standard Lagrange multiplier method (Bishop & Nasrabadi, 2006), we can obtain the closed-form solution for $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ in Equation A.6 shown below; and by equating the gradient of Equation A.2 to zero, we can obtain the closed-form solution for $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ in Equation A.7 shown below.
$\alpha_{k}=\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=k)}{I\cdot J},$ (A.6)
$\pi_{mn}^{(j^{\prime})}=\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=m)\cdot\Psi(y^{(i,j;j^{\prime})},n)}{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=m)}\mathrm{,}$ (A.7)
where $\Psi(y^{(i,j;j^{\prime})})=[\phi_{l}\cdot\mathbb{I}(y^{(i,j;j^{\prime})}_{l}=n)+\phi_{u}\cdot\mathbb{I}(y^{(i,j;j^{\prime})}_{u}=n]$, and $\mathbb{I}(\cdot)$ is an indicator function that takes the value $1$ when the internal declaration is true, and 0 otherwise.
A.2 More Experiment Setup
Table A.1: The models used in our experiments and their corresponding URLs.
The models used in the experiments are shown in Table n the following, we provide further descriptions of the four datasets used.
TriviaQA (Dubois et al., 2023). The TriviaQA dataset consists of 950K question-answer pairs from 662K documents on Wikipedia and the web. It is designed for complex question answering, where answers cannot always be directly obtained through span-based methods. The dataset includes both human-verified and machine-generated pairs. We utilize the dataset from the Smoothie (Guha et al., 2024), containing 1000 randomly selected samples from the original dataset.
GSM8K (Chen et al., 2021a). GSM8K (Grade School Math 8K) is a dataset of 8.5K math word problems designed for grade school students, with 7.5K training and 1K test problems. Each problem requires 2 to 8 steps to solve, focusing on basic arithmetic operations. The natural language solutions allow for evaluating models’ reasoning abilities. Given the high resource cost associated with using the full test set, we adopt existing approaches and utilize the publicly available dataset in Hu et al. (2024), which consist of 400 randomly selected samples from the original full test set.
MATH (Hendrycks et al., 2021). The MATH dataset consists of 12,500 challenging, competition-level math problems across topics like algebra, geometry, probability, and number theory. Each problem is accompanied by a detailed, step-by-step solution, making it ideal for training models to not only find answers but also generate reasoning and logical explanations. This dataset is highly valuable for advancing and evaluating models in solving complex mathematical problems. Also, given the high resource cost associated with using the full test set, we adopt existing approaches and utilize the publicly available dataset in Hu et al. (2024), which consist of 400 randomly selected samples from the original full test set.
AlpacaEval (Dubois et al., 2023). AlpacaEval consists of 805 instructions (Hu et al., 2024), including 252 from the self-instruct test set (Wang et al., 2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al., 2023), 80 from the Vicuna test set (Chiang et al., 2023), and 156 from the Koala test set (Vu et al., 2023).
A.3 More Experimental Results
In this section, we provide further experimental results, including: more variant results, win-tie-loss charts for models across all four datasets, and four case studies, each corresponding to one dataset.
[…]
Single LLM Llama-3.1-8B-Instruct 75.3 79.3 52.3 7.3 53.5
Mistral-7B-Instruct 72.7 64.3 26.5 10.4 43.5
Qwen2-7B-Instruct 63.0 88.5 59.8 15.2 56.6
Qwen2.5-7B-Instruct 62.5 91.5 69.3 27.6 62.7
Theoretical average 68.4 80.9 51.9 15.1 54.1
LLM Ensemble Random 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global 63.0 91.5 59.8 27.6 60.5
Smoothie-Local 73.6 85.5 61.8 18.3 59.8
Agent-Forest 70.5 86.8 61.0 22.1 60.1
LLM-PeerReview-Average 76.9 $\pm$ 0.1 92.7 $\pm$ 0.3 69.5 $\pm$ 0.2 30.4 $\pm$ 0.1 67.4
LLM-PeerReview-Weighted 77.0 $\pm$ 0.1 93.0 $\pm$ 0.2 71.0 $\pm$ 0.2 30.2 $\pm$ 0.1 67.8
Our variants (flipped-triple) Llama-3-8B-Selection 76.5 $\pm$ 0.2 90.8 $\pm$ 0.6 68.8 $\pm$ 0.5 29.6 $\pm$ 0.3 66.4
Mistral-7B-Selection 75.6 $\pm$ 0.3 90.8 $\pm$ 0.1 66.4 $\pm$ 0.3 25.9 $\pm$ 0.4 64.7
Qwen2-7B-Selection 74.2 $\pm$ 0.2 88.8 $\pm$ 0.6 61.7 $\pm$ 0.7 23.7 $\pm$ 0.3 62.1
Qwen2.5-7B-Selection 75.5 $\pm$ 0.2 92.1 $\pm$ 0.4 66.2 $\pm$ 0.6 28.1 $\pm$ 0.1 65.5
Our variants (single) Llama-3-8B-Selection 69.8 $\pm$ 0.3 83.7 $\pm$ 1.3 56.8 $\pm$ 0.4 21.5 $\pm$ 0.6 58.0
Mistral-7B-Selection 71.1 $\pm$ 0.9 82.9 $\pm$ 0.5 57.3 $\pm$ 0.5 18.6 $\pm$ 0.1 57.5
Qwen2-7B-Selection 70.9 $\pm$ 0.5 81.7 $\pm$ 0.6 53.4 $\pm$ 0.2 16.9 $\pm$ 0.3 55.7
Qwen2.5-7B-Selection 71.0 $\pm$ 0.8 83.2 $\pm$ 0.9 55.4 $\pm$ 0.5 23.8 $\pm$ 0.2 58.4
Table A.2: Main results (%).
Figure A.1: Win-tie-loss charts on four datasets.
More variant results.
By examining Table A.2, we observe the following: (i) The four variants of “single” obtained through the “naive point-wise scoring” method outperform the “theoretical average” across all four datasets. This suggests that these variant methods have a distinct advantage over random selection from model responses, implying they are useful; (ii) However, these four variants are still insufficiently useful when compared to the flipped-triple variant methods in Table A.3, indicating a performance disadvantage. By contrasting the performance of “variants (single)” and “variants (flipped-triple)”, it becomes clear that our proposed flipped-triple scoring strategy is indeed effective. This conclusion supports the assertion that the flipped-triple strategy, as designed in Section 2.1 of the main paper, mitigates the bias introduced by the naive point-wise scoring method. Furthermore, it is precisely upon this superior “flipped-triple” scoring strategy that the performance of our two primary variants, LLM-PeerReview-Average and LLM-PeerReview-Weighted, exhibits its ultimate superiority.
Win-tie-loss charts.
As a supplement to Figure 3 in the main text, we provide Figure A.1 here, showing the win-tie-loss charts for models across the four datasets. (i) Overall, on the TriviaQA dataset, LLAMA outperforms the other three models; on the GSM8K, MATH, and AlpacaEval datasets, Qwen2.5 consistently performs the best, although its advantage varies in strength across datasets; (ii) However, consistent with the analysis in Section 3.2 of the main text, observing the win-tie-loss charts across the four datasets reveals that, on a specific dataset, despite a model having a significant overall performance advantage, its performance on individual queries (i.e., samples) may not be better than a model with an overall suboptimal performance. This intuitive reality suggests that attempting LLM Ensemble makes a lot of sense, and a sufficiently sophisticated method could lead to even higher performance than the best-performing individual model.
Case studies.
Subsequently, in Figures A.2, A.3, A.4, and A.5, we present a case study for each corresponding dataset. Each case study illustrates the core computational process and final results of our method, as well as those of the comparison methods, Agent-Forest (Li et al., 2024c) and Smoothie (Guha et al., 2024), highlighting the key computational steps and outcomes. We provide a more detailed analysis of the first two datasets, as shown in the captions; the patterns observed in the remaining two datasets are consistent with those in the first two.
A.4 Prompts
In this section, we first introduce the prompts for each task across the corresponding datasets. We then introduce the scoring prompts, which are unique to our method in comparison to the baseline, designed specifically for scoring responses using the LLM-as-a-Judge technique.
Figure A.2: Case study on dataset TriviaQA. Analysis: For this query, the correct answer is “Leander Club”. For each response generation, we followed the approach in  Guha et al. (2024) and applied truncation techniques under the same configuration. (i) For our method, the results of our variant, LLM-PeerReview-Average, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [2.39, 2.39, 1.18, 1.18]; (ii) For the baseline Agent-Forest, it focuses on “which response has the highest cumulative BLEU similarity to all other responses”. Since Responses 1 and 2 are correct, while Responses 3 and 4 are incorrect, the cumulative BLEU similarity of Response 4 with all others is the highest, so Agent-Forest incorrectly selects Response 4; (iii) For the baseline Smoothie, its two variants do not consider the four responses for the current query during the calculation. Instead, they either focus on identifying the best overall model or the responses from the most similar neighboring query. As a result, both variants of Smoothie also make errors.
Figure A.3: Case study on dataset GSM8K. The ellipses “…” in the Responses represent the omission of less important content. Analysis: The correct answer for this query is “2”. (i) For our method, the results of our variant, LLM-PeerReview-Average, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [1.12, 1.00, 2.01, 2.90]; (ii) For the Agent-Forest method, it focuses on determining “which response has the highest cumulative BLEU similarity to all other responses”. Despite Response 2 not providing the correct final answer, Agent-Forest selects it because the weight (i.e., cumulative BLEU similarity score) is the highest. This is likely due to the high similarity between many words in the response compared to other responses. In this case, the simple analysis of “which response has the highest cumulative BLEU similarity” fails; (iii) For the Smoothie method, its two variants do not consider the four responses for the current query during the computation. Instead, they either focus on identifying the best overall model or consider the four responses of the most similar neighboring query. In contrast to the results shown in Figure A.2, here both variants of Smoothie obtained the correct result, despite the computation process being flawed.
Figure A.4: Case study on dataset MATH. For this query, the correct answer is “2500”. The ellipses “…” in the Responses represent the omission of less important content.
Figure A.5: Case study on dataset AlpacaEval. The ellipses “…” in the Responses represent the omission of less important content. For this query, the best response is Response 4, and the reference answer is “A word that represents people reacting to unpleasant events is ”resilience.” Resilience refers to the ability of individuals to cope with, adapt to, and recover from stress, adversity, trauma, or tragedy. It implies a form of mental toughness and flexibility that allows people to face challenges and bounce back from them.”
Prompt template for dataset TriviaQA. Answer the question below. Question: [placeholder of question] Answer:
Table A.4: Prompt template for dataset TriviaQA.
A prompt example (after query instantiation) for dataset TriviaQA. Answer the question below. Question: What amateur rowing club based at Henley on Thames, England, was founded in 1818? Answer:
Table A.4: A prompt example (after query instantiation) for dataset TriviaQA.
Prompt template for dataset GSM8K. Can you solve the following math problem? [placeholder of question]. Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.
Table A.5: Prompt template for dataset GSM8K.
A prompt example (after query instantiation) for dataset GSM8K. Can you solve the following math problem? A wall mural has four different colors of paint in it: red, white, purple, and yellow. There are equal amounts of red, white, and purple paint in the mural. Half the mural is yellow. If the mural used 12 pints of paint in all, how many pints of red paint was used? Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.
Table A.6: A prompt example (after query instantiation) for dataset GSM8K.
Prompt template for dataset MATH. Can you solve the following math problem? [placeholder of question]. Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.
Table A.7: Prompt template for dataset MATH.
A prompt example (after query instantiation) for dataset MATH. Can you solve the following math problem? If $c$ is a constant such that $x^{2}+100x+c$ is equal to the square of a binomial, then what is $c$? Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.
Table A.8: A prompt example (after query instantiation) for dataset MATH.
A prompt example (after query instantiation) for dataset AlpacaEval. Find a word that represents people reacting to unpleasant events.
Table A.9: A prompt example (after query instantiation) for dataset AlpacaEval.
Scoring prompt template for dataset TriviaQA. You are an expert evaluator tasked with assessing qualities of the three Responses generated for a Question/Instruction from the TriviaQA (factual recall) dataset in NLP. TriviaQA (factual recall) is a dataset designed for factual recall, containing a large number of question-answer pairs based on multiple sources such as Wikipedia and books, covering a wide range of topics including history, geography, culture, entertainment, science, and more. **Its purpose is to test how well models provide the correct facts/concepts when given factual questions.** You will be provided with a Question/Instruction and three Responses generated from different models. **Your task is to evaluate the qualities of the three Responses by focusing on the following evaluation criteria: For each Response, regardless of whether the entire Response contains irrelevant or nonsensical content, focus solely on whether the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question is mentioned/appears anywhere in the entire Response**: Simply search for whether the fact/concept that can correctly answer the Question is mentioned/appears anywhere in the entire Response. If it does, the Response is considered correct. For example, the fact “McDonnell Douglas” is correct for answering the Question “First flown on the 27th July 1972, who built the F15 Eagle fighter?”. In this case, regardless of whether the entire Response contains irrelevant or nonsensical content, you should only check for the mention of the correct fact/concept in various possible expressions (e.g., “McDonnell Douglas”, “mcdonell douglas”, “mcdonnell douglas aerospace”, “mcdonnell douglas corp”, “mc donell douglas”, etc.). If any of these variations are present in the Response, it is considered correct. In line with the above evaluation criteria, here are the scoring guidelines between [1, 2, 3, 4, 5]. Please provide **three precise integer scores between [1, 2, 3, 4, 5]** by **strictly and meticulously adhering to the following scoring guidelines**: 1 = Regardless of whether the entire Response contains irrelevant or nonsensical content, the Response **does not contain** the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question. 2 = Regardless of whether the entire Response contains irrelevant or nonsensical content, the Response **most likely does not contain** the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question. 3 = Regardless of whether the entire Response contains irrelevant or nonsensical content, the Response **seems likely (20%-60%) to contain** the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question. 4 = Regardless of whether the entire Response contains irrelevant or nonsensical content, the Response **is highly likely (60%-99%) to contain** the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question. 5 = Regardless of whether the entire Response contains irrelevant or nonsensical content, the Response **100% contains** the fact/concept (including different expressions referring to the same fact/concept, such as variations in capitalization, aliases, or minor formatting issues) that can correctly answer the Question. Note: 1) **Output format: Please return only a Python dictionary in the format “Score for the Response One”: x, “Score for the Response Two”: y, “Score for the Response Three”: z where x,y,z are the placeholders that you should replace with the integer scores for each Response. Do NOT include any additional opening, closing, explanation, or formatting.** 2) **Objectivity**: - **Do not let the order of the Responses to introduce any bias in your scoring**: As an expert evaluator, you should strictly follow the scoring guidelines, ensuring that the order of the Responses does not influence your judgment in any way. You can reconsider and evaluate the Responses multiple times internally to absolutely ensure that “order” does not affect your final scores. - **Please avoid any bias: Do NOT let your judgment be swayed by any conservatism or exaggeration. Carefully consider the content of the three Responses,and strictly adhere to the scoring guidelines to provide three accurate scores.** Now here are the Question/Instruction, the Responses, and your judge result: Question/Instruction: [placeholder of question] Response One: [placeholder of response1] Response Two: [placeholder of response2] Response Three: [placeholder of response3] Your Python dictionary containing the three scores:
Table A.10: Scoring prompt template for dataset TriviaQA.
Scoring prompt template for dataset GSM8K. You are an expert evaluator tasked with assessing the qualities of three Responses to math problems from the GSM8K dataset in NLP. You will be provided with an Instruction/Question and three Responses generated by different models. Your task is to evaluate the qualities of three Responses by focusing on the following evaluation criteria: After carefully analyzing the math problem and deriving a final answer yourself, for each Response, regardless of whether the mathematical reasoning or analysis in the Response is correct or reasonable, or whether the Response contains irrelevant, redundant or repetitive content, your evaluation should focus solely on two possible scenarios (**for each Response, only one of the following two cases applies**): **(Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: Simply search for whether the correct answer to the math Question is mentioned/appears in any of these sentences.** **(Case B) If the Response does not contain any sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: Simply check whether **the LAST single numerical number in the entire Response** is the correct answer (usually a single numerical value) to the math Question.** In line with the above evaluation criteria, here are the scoring guidelines between [1, 2, 3]. Please provide **three precise integer scores between [1, 2, 3]** by **strictly and meticulously adhering to the following scoring guidelines**: 1 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **do not contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **is not** the correct answer (usually a single numerical value) to the math Question. 2 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **seems likely or very likely to contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **seems likely or very likely to be** the correct answer (usually a single numerical value) to the math Question. 3 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **100% contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **100% is** the correct answer (usually a single numerical value) to the math Question. Note: 1) **Output format: Please return only a Python dictionary in the format {“Score for the Response One”: x, “Score for the Response Two”: y, “Score for the Response Three”: z} where x,y,z are the placeholders that you should replace with the integer scores for each Response. Do NOT include any additional opening, closing, explanation, or formatting.** 2) **Objectivity**: - **Do not let the order of the Responses to introduce any bias in your scoring**: As an expert evaluator, you should strictly follow the scoring guidelines, ensuring that the order of the Responses does not influence your judgment in any way. You can reconsider and evaluate the Responses multiple times internally to absolutely ensure that “order” does not affect your final scores. - **Please avoid any bias: Do NOT let your judgment be swayed by any conservatism or exaggeration. Carefully consider the content of the three Responses,and strictly adhere to the scoring guidelines to provide three accurate scores.** Now here are the Instruction/Question, the Responses, and your judge result: Instruction/Question: [placeholder of question] Response One: [placeholder of response 1] Response Two: [placeholder of response 2] Response Three: [placeholder of response 3] Your Python dictionary containing the three scores:
Table A.11: Scoring prompt template for dataset GSM8K.
Scoring prompt template for dataset MATH. You are an expert evaluator tasked with assessing the qualities of three Responses to math problems from the MATH dataset in NLP. You will be provided with an Instruction/Question and three Responses generated by different models. Your task is to evaluate the qualities of three Responses by focusing on the following evaluation criteria: After carefully analyzing the math problem and deriving a final answer yourself, for each Response, regardless of whether the mathematical reasoning or analysis in the Response is correct or reasonable, or whether the Response contains irrelevant, redundant or repetitive content, your evaluation should focus solely on two possible scenarios (**for each Response, only one of the following two cases applies**): **(Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: Simply search for whether the correct answer to the math Question is mentioned/appears in any of these sentences.** **(Case B) If the Response does not contain any sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: Simply check whether **the LAST single numerical number in the entire Response** is the correct answer (usually a single numerical value) to the math Question.** In line with the above evaluation criteria, here are the scoring guidelines between [1, 2, 3]. Please provide **three precise integer scores between [1, 2, 3]** by **strictly and meticulously adhering to the following scoring guidelines**: 1 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **do not contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **is not** the correct answer (usually a single numerical value) to the math Question. 2 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **seems likely or very likely to contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **seems likely or very likely to be** the correct answer (usually a single numerical value) to the math Question. 3 = Regardless of how irrelevant, unreasonable, or unformatted other content in the Response is: (Case A) If the Response contains a sentence/sentences containing “the answer is”, “The answer is”, “answer”, or “Answer”: These “answer” sentences **100% contain** the correct answer (usually a single numerical value) to the math Question. (Case B) If the Response does not contain any sentence containing “the answer is”, “The answer is”, “answer”, or “Answer”: **The last numerical value** that appears in the entire Response **100% is** the correct answer (usually a single numerical value) to the math Question. Note: 1) **Output format: Please return only a Python dictionary in the format {“Score for the Response One”: x, “Score for the Response Two”: y, “Score for the Response Three”: z} where x,y,z are the placeholders that you should replace with the integer scores for each Response. Do NOT include any additional opening, closing, explanation, or formatting.** 2) **Objectivity**: - **Do not let the order of the Responses to introduce any bias in your scoring**: As an expert evaluator, you should strictly follow the scoring guidelines, ensuring that the order of the Responses does not influence your judgment in any way. You can reconsider and evaluate the Responses multiple times internally to absolutely ensure that “order” does not affect your final scores. - **Please avoid any bias: Do NOT let your judgment be swayed by any conservatism or exaggeration. Carefully consider the content of the three Responses,and strictly adhere to the scoring guidelines to provide three accurate scores.** Now here are the Instruction/Question, the Responses, and your judge result: Instruction/Question: [placeholder of question] Response One: [placeholder of response 1] Response Two: [placeholder of response 2] Response Three: [placeholder of response 3] Your Python dictionary containing the three scores:
Table A.12: Scoring prompt template for dataset MATH.
Scoring prompt template for dataset AlpacaEval (Part 1). You are an expert evaluator tasked with assessing the qualities of three Responses generated for an Instruction from the well-known instruction-following dataset AlpacaEval in NLP. You will be provided with an Instruction and three Responses generated by different models. Your task is to evaluate the overall qualities of three Responses by focusing on the following evaluation criteria: Always keep in mind the “Core Evaluation Criteria” — “Evaluate whether the Response effectively addresses the Instruction from a human perspective.” Also, please focus on the following **FOUR Specific Evaluation Criteria**: **Basic Requirement: Understanding the Question & Response Relevance**: Does the Response understand and follow the Instruction without going off-topic or irrelevant? **Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness**: Does the Response correctly, reasonably, and logically answer the Instruction? Overall, does the Response meet the core requirements of the Instruction in terms of usefulness and effectiveness? **Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value**: In cases where needed (especially for more open-ended Instructions), does the Response thoroughly cover all aspects of the Instruction’s requirements, including all key points, relevant aspects, necessary information, steps for problem-solving needed to perfectly answer the Instruction? Does the Response go beyond surface-level answers and provide valuable, useful, creative, more in-depth information, offering added value and making the response more perfect? **Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding**: Does the Response use appropriate language that is clear, natural, coherent, fluent, concise, and easy to understand? Does the Response avoid vague, unnecessarily verbose, or repetitive statements? In line with the above evaluation criteria, here are the scoring guidelines between [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Please provide **three precise integer scores between [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]** by **strictly and meticulously adhering to the following scoring guidelines**: **1 = From a human perspective, overall, the Response “fails to address the Instruction’s needs”, and it is a “Unrelevant, Unacceptable Response”.** Because: It fails to meet the above “Basic Requirement: Understanding the Question & Response Relevance”, let alone the more advanced requirements. **2 = From a human perspective, overall, the Response “partially addresses the Instruction’s needs to a small extent”, and it is a “Relevant, Low-Utility Response”.** Because: It meets the above “Basic Requirement: Understanding the Question & Response Relevance”, but **“has too many serious and obvious shortcomings in some of the the following aspects”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness”, “Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **3 = From a human perspective, overall, the Response “partially addresses the Instruction’s needs to a moderate extent”, and it is a “Relevant, Moderately Useful Response”.** Because: It meets the above “Basic Requirement: Understanding the Question & Response Relevance”, but **“has several serious and obvious shortcomings in some of the the following aspects”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness”, “Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **4 = From a human perspective, overall, the Response “partially addresses the Instruction’s needs to a large extent”, and it is a “Relevant, Useful Response”.** Because: It meets the above “Basic Requirement: Understanding the Question & Response Relevance”, but **“has several obvious shortcomings in some of the the following aspects”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness”, “Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **5 = From a human perspective, overall, the Response “almost addresses the Instruction’s needs”, and it is a “Relevant, Useful, Minimally Acceptable Response”.** Because: It meets the above “Basic Requirement: Understanding the Question & Response Relevance”, but **“has noticeable shortcomings in some of the the following aspects”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness”, “Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **6 = From a human perspective, overall, the Response “sufficiently addresses the Instruction’s needs”, and it is a “Relevant, Useful, Acceptable Response.** Because: It meets the above “Basic Requirement: Understanding the Question & Response Relevance”, and it **“has minor, subtle shortcomings in some of the the following aspects”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and Overall Usefulness/Effectiveness”, “Content-Related Advanced Requirement: Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Language-Related Advanced Requirement: Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **7 = From a human perspective, overall, the Response “effectively addresses the Instruction’s needs”, and it is a “Relevant, Useful, Acceptable, Good Response”.** Because: It meets the “Basic Requirement: Understanding the Question & Response Relevance”, and **“there may be minor room for improvement in the following criteria”: “Core Requirement: Correctness, Reasonableness, Logical Consistency, and overall Usefulness/Effectiveness”, “Advanced Requirement One (Content-Related): Completeness, Thoroughness, Depth, Creativity, and Added Value”, “Advanced Requirement Two (Language-Related): Clarity, Naturalness, Coherence, Fluency, Conciseness, and Ease of Understanding”**. **8 = From a human perspective, overall, the Response “excellently addresses the Instruction’s needs”, and it is a “Relevant, Useful, Strongly Acceptable, Excellent Response”.** Because: It excels in all “FOUR Specific Evaluation Criteria”, and there seems to be no room for improvement. **9 = From a human perspective, overall, the Response “perfectly addresses the Instruction’s needs“, and it is a “Relevant, Useful, Strongly Acceptable, Perfect Response”.** Because: It fully excels in all above FOUR Specific Evaluation Criteria, with NO room for improvement. **10 = From a human perspective, overall, the Response “exceeds expectations and very perfectly addresses” the Instruction’s needs, and it is a “Relevant, Useful, Very Strong Acceptable, Very Perfect Response”.** Because: It fully excels in all above FOUR Specific Evaluation Criteria, with NO room for improvement.
Table A.13: Scoring prompt template for dataset AlpacaEval (Part 1).
Scoring prompt template for dataset AlpacaEval (Part 2). Note: 1) **Output format: Please return only a Python Dictionary, 100% STRICTLY following in the format “Score for the Response One”: x, “Score for the Response Two”: y, “Score for the Response Three”: z where x,y,z are the placeholders that you should replace with the integer scores for each Response. NOTE: You are NOT allowed to modify the “key” of the Dictionary under any circumstances. Do NOT include any additional opening, closing, explanation, or formatting.** 2) **Objectivity**: - **Do not let the order of the Responses to introduce any bias in your scoring**: As an expert evaluator, you should strictly follow the scoring guidelines, ensuring that the order of the Responses does not influence your judgment in any way. You can reconsider and evaluate the Responses multiple times internally to absolutely ensure that “order” does not affect your final scores. - **Please avoid any bias: Do NOT let your judgment be swayed by any conservatism or exaggeration. Do NOT allow the length of the responses to influence your evaluation. Carefully consider the content of the three Responses, and strictly adhere to the scoring guidelines to provide three precise scores.** Now here are the Instruction, the Responses, and your judge result: Instruction: [placeholder of question] Response One: [placeholder of response1] Response Two: [placeholder of response2] Response Three: [placeholder of response3] Your Python dictionary containing the three scores:
Table A.14: Scoring prompt template for dataset AlpacaEval (Part 2).